OpenStudy (anonymous):

Fan & Medals!! Plz help (3/2)x = (-1/2)x + 4 -- And solve for x

OpenStudy (anonymous):

do you have choose

OpenStudy (anonymous):

Add 1/2x to both sides: \(\dfrac{3}{2}x + \dfrac{1}{2}x = 4\) Add the two fractions: \(\dfrac{3 + 1}{2}x = 4\) Can you add 3 + 1?

OpenStudy (anonymous):

3/2x = -1/2x + 4 3x = -x + 8 4x = 8 x = 2

OpenStudy (anonymous):

@BethLM39 Don't give answers.

OpenStudy (anonymous):

I agree with @iGreen.

OpenStudy (anonymous):

Don't tell me what to do.

OpenStudy (anonymous):

I should probably tell you not to do that, because otherwise you will be banned..read the rules. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I don't think @BethLM39 meant to put the answer and if he did than there is nothing anyone can really do about.

OpenStudy (anonymous):

I think @BethLM39 is a she..lol.

OpenStudy (anonymous):

"I will encourage and guide." Which is what I did. I wrote out the steps to show how the problem is completed. I could've just said "x=2" and been on my merry way.

OpenStudy (anonymous):

whatever still you should not be telling other people what to do.

OpenStudy (anonymous):

it is their choose to put the answer down than get banned. let them make that choose.

OpenStudy (anonymous):

we should be helping @LOLAri and not fighting about the rules

OpenStudy (anonymous):

its ok guys btw can you guys look at my new question just posted it=D

OpenStudy (anonymous):

yes